OpenStudy (anonymous):

Use the quadratic formula to solve for x 3x^2+3x-7=0 (If there's more than one solution, separate them with commas)

OpenStudy (anonymous):

\[(-3-\sqrt{93})\div6,(-3+\sqrt{93})/6\]

OpenStudy (agreene):

Vood is correct, but to make it a bit easier to read: \[\frac16(-3\pm\sqrt{93})\]